OpenStudy (darkbluechocobo):

Help with true and false questions about cosecant and secant functions

OpenStudy (darkbluechocobo):

1.The vertical asymptotes of the cosecant function are located at the same values of x for which the sine function is equal to 0. 2.The vertical asymptotes of the secant function are located at the same values of x for which the sine function is equal to 0.

OpenStudy (freckles):

secant=1/cosine so when cosine is 0 secant is undefined right?

OpenStudy (darkbluechocobo):

Yeh I just saw this in my notes so it would be the secant that is true and false for cosecant?

OpenStudy (darkbluechocobo):

@freckles

OpenStudy (freckles):

wait which are you saying is true?

OpenStudy (freckles):

1 or 2?

OpenStudy (darkbluechocobo):

so im saying 1 is false and 2 is true

OpenStudy (freckles):

why would you say 2 is true when I just said that secant=1/cosine and when cosine is 0 secant is undefined?

OpenStudy (freckles):

also we will look at 1 again after we have finished looking at 2

OpenStudy (darkbluechocobo):

I miss thought is all then s: see I just got done dealing with Cotangent and if Cos=0 that means that it would be true but we arent dealing with that

OpenStudy (freckles):

\[\sec(x)=\frac{1}{\cos(x)} \\ \text{ if } \cos(x)=0 \text{ then } \sec(x)=\frac{1}{0} \\ \text{ you cannot divide by 0 }\]

OpenStudy (darkbluechocobo):

Yes that is true you cannot

OpenStudy (freckles):

so that means sec(x) has a vertical asymptote when cos(x) is 0

OpenStudy (freckles):

so number 2 is true or false?

OpenStudy (darkbluechocobo):

I am confused s: if cos is 0 that means it is undefined

OpenStudy (freckles):

exactly

OpenStudy (darkbluechocobo):

so how does that work?

OpenStudy (darkbluechocobo):

So if it is undefined it has vertical asymptotes?

OpenStudy (darkbluechocobo):

oh wait

OpenStudy (freckles):

usually 0/0 means you have a hole but if you have a/0 where a isn't 0 then you definitely have a vertical asymptote Remember f(x)=1/x That function looks like:  1/x is undefined when x=0 so since we have 1/0 we definitely have a vertical asymptote at x=0

OpenStudy (darkbluechocobo):

2 is false because it says sine not cosine

OpenStudy (freckles):

right if sin=0 then cos is definitely not zero at the same time

OpenStudy (freckles):

now what about number 1?

OpenStudy (darkbluechocobo):

Yes this is true aha sorry

OpenStudy (darkbluechocobo):

I would say that it\[\cos(x)=\frac{ 1 }{ \sec(x) }\]

OpenStudy (freckles):

1.The vertical asymptotes of the cosecant function are located at the same values of x for which the sine function is equal to 0. remember cosecant=csc csc(x)=1/sin(x) correct?

OpenStudy (darkbluechocobo):

Omg

OpenStudy (darkbluechocobo):

*FACEPALMS*

OpenStudy (darkbluechocobo):

I was like cosecant must be COS *FACEPALMS*

OpenStudy (freckles):

cosine=cos sine=sin cosecant=csc secant=sec cotangent=cot tangent=tan

OpenStudy (darkbluechocobo):

Yes yes I failed sorry my mind just went all stupid mode lool

OpenStudy (freckles):

lol anyways this is basically the same question before except with a different story ending